# Kernel Three Pass Regression Filter

### v3

@@ 3. Run time series regression of $y_{t+h}$ on predictive factors $\hat{\boldsymbol{F}}_{t}$,
[…]
Figure 2: Implementation of the Kernel Three Pass regression filter for the case T=4 and L=3 relevant factors. The variables $z^{(1)}\dots z^{(3)}$ and $f^{(1)}\dots f^{(3)}$ are the vectors representing the time series of the respective variables. $X_{s}$, (resp $\varphi(X_{s}))$ represents the cross section of $\boldsymbol{X}$ (resp $\varphi(\boldsymbol{X})$) in period $s$.
@@
$=\mathbf{\iota}_{T}\bar{y}+\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\boldsymbol{S}_{\varphi(\boldsymbol{X})Z}\left(\boldsymbol{S}_{\varphi(\boldsymbol{X})Z}^{\prime}\boldsymbol{S}_{\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})}\boldsymbol{S}_{\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})}\right)^{-1}\boldsymbol{S}_{\varphi(\boldsymbol{X})Z}^{\prime}\boldsymbol{S}_{\varphi(\boldsymbol{X})y}$
@@
These expressions are obtained by simply replacing $\boldsymbol{X}$ by $\varphi(\boldsymbol{X})$ in the three-pass regression filter of Kelly & Pruitt, (2015). As evident from the expression of $\hat{\boldsymbol{F}}^{\prime}$, the filtration process applied on the transformed predictor space results in a favorable scenario where the eventual estimate of the factor(s) depends upon the transformed predictors only through their dot products in the transformed space. This holds true for $\hat{\boldsymbol{\beta}}$ and $\hat{\boldsymbol{y}}$ as well.
@@
The Kernel three-pass regression, like the linear 3PRF, relies on the availability of suitable proxies. Kelly & Pruitt, (2015) show that such proxies can always be constructed using the target variable $\boldsymbol{y}$. That process is explained in table-3 below.
0. Initialize $\boldsymbol{r}_{0}=\boldsymbol{y}$. For $k=1,\ldots,L$. (L is the total number of proxies)
@@
2. Compute the k3PRF for target $\boldsymbol{y}$ using cross-section $\boldsymbol{X}$ using statistical proxies 1 through $k$.
Denote the resulting forecast $\hat{\boldsymbol{y}}_{k}$.
3. Calculate $\boldsymbol{r}_{k}=\boldsymbol{y}-\hat{\boldsymbol{y}}_{k}$, advance $k$, and go to step 1.
@@
    $\mathbb{E}\left\|\phi_{i}\right\|^{4}\leq R$,$M^{-1}\sum_{j=1}^{M}\phi_{j}\underset{N\rightarrow\infty}{\stackrel{{\scriptstyle p}}{{\longrightarrow}}}\boldsymbol{0}$, $M^{1/2}\left(\dfrac{\boldsymbol{\Phi}^{\prime}\boldsymbol{\Phi}}{M}-\mathcal{P}\right)=\boldsymbol{O}_{p}(1)$.
@@
Assumption 2.1 requires that our factors are regular in the sense that their covariance matrix is well-behaved asymptotically. Assumption 2.2 is an adaptation from Kelly & Pruitt, (2015). Since we assume a factor structure on the transformed space instead of the original predictor space, the normalization in various terms features $M$ and not $N$, where $M$ is the dimension of our transformed space. Assumptions 2.3-2.5, borrowed from Kelly & Pruitt, (2015), impose regularity on various error processes.
@@
Assumption 3.1-3.2 allow various forms of weak cross-sectional and temporal dependence between the idiosyncratic components of the transformed predictors. These assumptions characterize our ‘Approximate’ factor model. The terminology of approximate, as opposed to a strict factor model, alludes to the allowance of these weak correlations, as outlined by Chamberlain & Rothschild, (1983). These assumptions are standard in the literature; see Bai, (2003). Assumption 3.4-3.7 are either borrowed from or are weaker versions of Assumptions in Kelly & Pruitt, (2015). They are reasonable because each of them involves a product of orthogonal series.
@@
Assumption 4 is a normalization assumption that is common in factor model literature. It pertains to the non-identifiability of the true factor(s). It is well known that only the vector space spanned by the factor(s) can be consistently estimated but not the factor themselves. Imposing some normalization condition for the uniqueness of solution(s) is common in literature.
@@
Assumption 5 outlines the utility of using proxies. Proxies are target-relevant in the sense that they only load on the factor(s) that have any explanatory power for the target. Non-singularity of $\boldsymbol{\Lambda}_{f}$ ensures that none of the proxies are redundant.
4 Results
@@
Define $\delta_{MT}\equiv min\{\sqrt{M},\sqrt{T}\}$. Define $\boldsymbol{H}_{f}\equiv\hat{\boldsymbol{F}}_{A}\hat{\boldsymbol{F}}_{B}^{-1}\mathbf{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}$ where, $\hat{\boldsymbol{F}}_{A}=T^{-1}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}$ and $\hat{\boldsymbol{F}}_{B}=M^{-1}T^{-2}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X}^{\prime})\boldsymbol{J}_{T}\boldsymbol{Z}$
@@
This theorem establishes the estimated factor(s) convergence to the true factors up to a rotation. It is well known in the literature on factor models [^5] This feature of inherent unidentifiability has been emphasized in Bai, (2003), Kelly & Pruitt, (2015) among other papers. The normalization imposed in assumption 5 is done to handle this issue., that true underlying factor(s) are not identifiable; we instead estimate a rotated version of the true factors, which preserves their span.
Define $\boldsymbol{G}_{\beta}\equiv\hat{\boldsymbol{\beta}}_{1}^{-1}{\hat{\boldsymbol{\beta}}_{2}}\left[\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}\right]^{-1}\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}$, where
$\hat{\boldsymbol{\beta}}_{1}=\hat{\boldsymbol{F}}_{A}$ and $\hat{\boldsymbol{\beta}}_{2}=\hat{\boldsymbol{F}}_{B}$
@@
$\hat{\boldsymbol{\beta}}-\boldsymbol{G}_{\beta}\boldsymbol{\beta}=\boldsymbol{O}_{p}(\delta_{MT}^{-1}).$
${\boldsymbol{H}_{f}}^{\prime}\boldsymbol{G}_{\beta}=\boldsymbol{I}$
This theorem establishes the convergence of the predictive coefficients to a rotation of the true coefficients. Just like in the case of factor(s), true coefficients are not identifiable and we instead estimate their rotation. The rates established in Theorems 1 and 2 differ from the rates established in Kelly & Pruitt, (2015) and the reason is that our definition of rotation matrices $\boldsymbol{H}_{f}$ and $\boldsymbol{G}_{\beta}$ are different from Kelly & Pruitt, (2015). (See Remark 1).
@@
As highlighted in Bai & Ng, (2006) and also emphasized in Kelly & Pruitt, (2015), the presence of matrices $\boldsymbol{H}_{f}$ and $\boldsymbol{G}_{\beta}$ in Theorem 1 and 2 highlight the fact we are essentially estimating a vector space. These theorems “pertain to the difference between $\left[\hat{\boldsymbol{F}}_{t}/\hat{\boldsymbol{\beta}}\right]$ and the space spanned by $\left[\boldsymbol{F}_{t}/\boldsymbol{\beta}\right]$”. The product $\boldsymbol{H}_{f}^{\prime}\boldsymbol{G}_{\beta}$ equals an identity matrix, cancelling the rotations in the estimated coefficients and the factors; thereby consistently estimating direction spanned by $\boldsymbol{\beta}^{\prime}\boldsymbol{F}_{t}$. However, this characteristic is absent in Theorems 5 and 6 of Kelly & Pruitt, (2015). The matrices $\boldsymbol{H}$ and $\boldsymbol{G}_{\beta}$ as defined in their paper do not necessarily yield a product that equals an identity matrix.
@@
$\hat{y}_{t+h}-\mathbb{E}_{t}y_{t+h}=O_{p}(\delta_{MT}^{-1})$
Combining Theorem 1 and 2, the convergence $\hat{y}_{t+h}$ of follows directly. Our proof, unlike Kelly & Pruitt, (2015) uses the convergence results for the estimated factor(s) and coefficients to obtain this result.
@@
The rates established in Theorem 1, 2 and 3 are different from the result in Kelly & Pruitt, (2015) where the corresponding rates are $O_{p}({T}^{-1/2})$, $O_{p}({T}^{-1/2})$ and $O_{p}({N}^{-1/2})$ [^6] For our case, it should have been $O_{p}({M}^{-1/2})$ as per their theorem since we apply 3PRF to the transformed M-dimesnional space. respectively (see Theorems 4, 5 and 6 in their paper). For Theorem 1 and 2, the difference is explained by a different definition of the rotation matrices in our paper (see Remark 1). For establishing the convergence of $\hat{y}_{t+h}$, their proof follows two steps. First they show that $\hat{y}_{t+h}-\tilde{y}_{t+h}=O_{p}({T}^{-1/2})$, where $\tilde{y}_{t+h}$ is defined in their appendix. Then then they argue that $\sqrt{T}\tilde{y}_{t+h}\underset{T,N\rightarrow\infty}{\longrightarrow}\mathbb{E}_{t}y_{t+h}$. Since $\tilde{y}_{t+h}$ [^7] The definition of $\tilde{y}_{t+h}$ and fact that it is $O_{p}(1)$ can be seen from the proof of Theorem 6 of Kelly & Pruitt, (2015) is $O_{p}(1)$, $\sqrt{T}\tilde{y}_{t+h}$ would diverge to infinity and their statement would be false. We presume that they erroneously wrote this and instead wanted to imply that $\sqrt{T}\left(\tilde{y}_{t+h}-\mathbb{E}_{t}y_{t+h}\right)\underset{T,N\rightarrow\infty}{\longrightarrow}0$. However this statement is false because $\tilde{y}_{t+h}-\mathbb{E}_{t}y_{t+h}$ has random elements which converge to 0 at a rate which is $O_{p}(M^{-1/2})+O_{p}(T^{-1/2})=O_{p}(\delta_{MT}^{-1}).$

### v4

@@ 3. Run time series regression of $y_{t+h}$ on predictive factors $\hat{\boldsymbol{F}}_{t}$,
$z^{(1)}$$z^{(2)}$$z^{(3)}$$X_{1}$$X_{2}$$X_{3}$$X_{4}$$\varphi(X_{1})$$\varphi(X_{2})$$\varphi(X_{3})$$\varphi(X_{4})$$f^{(1)}$$f^{(2)}$$f^{(3)}$$y$
Figure 2: Implementation of the Three Pass regression filter for the case T=4 and L=3 relevant factors. The variables $z^{(1)}\dots z^{(3)}$ and $f^{(1)}\dots f^{(3)}$ are the vectors representing the time series of the respective variables. $X_{s}$, (resp $\varphi(X_{s}))$ represents the cross section of $\boldsymbol{X}$ (resp $\varphi(\boldsymbol{X})$) in period $s$.
@@
$=\mathbf{\iota}_{T}\bar{y}+\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\boldsymbol{S}_{\varphi(\boldsymbol{X})Z}\left(\boldsymbol{S}_{\varphi(\boldsymbol{X})Z}^{\prime}\boldsymbol{S}_{\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})}\boldsymbol{S}_{\varphi(\boldsymbol{X})\boldsymbol{Z}}\right)^{-1}\boldsymbol{S}_{\varphi(\boldsymbol{X})Z}^{\prime}\boldsymbol{S}_{\varphi(\boldsymbol{X})y}$
@@
These expressions are obtained by simply replacing $\boldsymbol{X}$ by $\varphi(\boldsymbol{X})$ in the three-pass regression filter of kelly. [^4] The only difference is that the matrix $\boldsymbol{J}_{N}$, used in kelly to demean along the cross-sectional dimension, is absent here. By omitting the intercept in the regression of $\boldsymbol{X}$, we avoid the need for this demeaning step. This is justified because we work with the kernel function rather than explicit transformed features, so the transformation is implicit and passes 1 and 2 are infeasible in practice. Such an approach is standard in kernel-based methods like kernel PCA, where demeaning is performed along the temporal (sample) dimension but not along the transformed predictor dimension. As evident from the expression of $\hat{\boldsymbol{F}}^{\prime}$, the filtration process applied on the transformed predictor space results in a favorable scenario where the eventual estimate of the factor(s) depends upon the transformed predictors only through their dot products in the transformed space. This holds true for $\hat{\boldsymbol{\beta}}$ and $\hat{\boldsymbol{y}}$ as well.
@@
The key insight is that the variation in the predictors that is relevant for forecasting $y_{t+h}$ may lie on a low-dimensional, nonlinear manifold in the original predictor space. After applying a kernel-induced transformation $\varphi(\cdot)$, this nonlinear structure becomes much closer to linear in the high-dimensional feature space spanned by the transformed features.
By assuming that $\varphi(x_{t})$ admits an approximate factor structure, we are effectively stating that the predictive manifold becomes well-approximated by a low-dimensional linear subspace in the transformed feature space. Therefore, while the conditional expectation $E[y_{t+h}\mid x_{t}]$ may be highly nonlinear in the original $N$-dimensional predictors, the same object becomes linear in a small number of relevant latent factors $f_{t}$ after the transformation $\varphi(\cdot)$.
This delivers the same idea as kelly: forecasting is driven by a few latent factors that summarize the relevant predictive variation. The only difference is that, rather than extracting these factors from the raw predictors, we extract them after a kernel-induced nonlinear transformation, which makes nonlinear predictive structure amenable to linear factor methods.
The Kernel three-pass regression, like the linear 3PRF, relies on the availability of suitable proxies. kelly show that such proxies can always be constructed using the target variable $\boldsymbol{y}$. That process is explained in Table-2.
0. Initialize $\boldsymbol{r}_{0}=\boldsymbol{y}$. For $k=1,\ldots,L$ (where $L$ is the total number of proxies).
@@
2. Compute the k3PRF forecast for target $\boldsymbol{y}$ using cross-section $\boldsymbol{X}$ and statistical proxies $1$ through $k$. Denote the resulting forecast by $\hat{\boldsymbol{y}}_{k}$.
3. Update the residual: $\boldsymbol{r}_{k}=\boldsymbol{y}-\hat{\boldsymbol{y}}_{k}$. Then increment $k$ and return to Step 1.
@@
    $\mathbb{E}\left\|\phi_{i}\right\|^{4}\leq R$$M^{-1}\sum_{j=1}^{M}\phi_{j}\underset{M\rightarrow\infty}{\stackrel{{\scriptstyle p}}{{\longrightarrow}}}\boldsymbol{0}$, $M^{1/2}\left(\dfrac{\boldsymbol{\Phi}^{\prime}\boldsymbol{\Phi}}{M}-\mathcal{P}\right)=\boldsymbol{O}_{p}(1)$.
@@
Assumption 2.1 imposes regularity on the latent factors by requiring their covariance matrix to converge to a well-behaved, non-stochastic limit, denoted $\boldsymbol{\Delta}_{F}$, a $K\times K$ matrix. Assumption 2.2, adapted from kelly, imposes an analogous condition on the factor loadings, treating $\mathcal{P}$ as a non-stochastic $K\times K$ matrix representing their limiting cross-sectional covariance. Likewise, $\boldsymbol{\Delta}_{\omega}$ is a non-stochastic $L\times L$ matrix capturing the population covariance of the proxy noise. Since we assume a factor structure on the transformed space instead of the original predictor space, the normalization in various terms features $M$ and not $N$, where $M$ is the dimension of our transformed space. Assumptions 2.3-2.5, borrowed from kelly, impose regularity on various error processes.
@@
Assumption 3.1-3.2 allow various forms of weak cross-sectional and temporal dependence between the idiosyncratic components of the transformed predictors. These assumptions characterize our ‘Approximate’ factor model. The terminology of approximate, as opposed to a strict factor model, alludes to the allowance of these weak correlations, as outlined by chamberlain. These assumptions are standard in the literature; see bai2003. Assumption 3.4-3.7 are either borrowed from or are weaker versions of Assumptions in kelly. They are reasonable because each of them involves a product of orthogonal series.
@@
Assumption 4 characterizes the covariance matrices defined in Assumption 2. This is a normalization assumption and is common in factor model literature, see kelly Assumption 5. It pertains to the non-identifiability of the true factor(s). It is well known that only the vector space spanned by the factor(s) can be consistently estimated but not the factor themselves. Imposing some normalization condition for the uniqueness of solution(s) is common in literature.
@@
Assumption 5 outlines the role and structure of the proxies in our framework. The proxy loading matrix is of the form $\boldsymbol{\Lambda}=[\boldsymbol{\Lambda}_{f}\;\;\mathbf{0}]$, so that the proxies load only on the relevant factors and have zero loading on the irrelevant factors ($\boldsymbol{\Lambda}_{g}=\mathbf{0}$). Non-singularity of $\boldsymbol{\Lambda}_{f}$ ensures that the proxies are linearly independent and none are redundant. Implicitly, this also requires that the number of proxies $L$ matches the number of relevant factors $K_{f}$, so that the relevant factor space is fully spanned by the proxies. In other words, the common component of the proxies captures all variation in the relevant factors, while none of the proxy variation is due to irrelevant factors.
4 Theoretical Results
@@
Define $\delta_{MT}\equiv min\{\sqrt{M},\sqrt{T}\}$. Define $\boldsymbol{H}_{f}\equiv\hat{\boldsymbol{F}}_{A}\hat{\boldsymbol{F}}_{B}^{-1}\mathbf{\Lambda}_{f}\boldsymbol{\Delta}_{f}$ where, $\hat{\boldsymbol{F}}_{A}=T^{-1}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}$ and $\hat{\boldsymbol{F}}_{B}=M^{-1}T^{-2}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X}^{\prime})\boldsymbol{J}_{T}\boldsymbol{Z}$, and $\boldsymbol{\Delta}_{f}$ denotes the submatrix of $\boldsymbol{\Delta}_{F}$ formed by its first $K_{f}$ rows and first $K_{f}$ columns.
@@
This theorem establishes the estimated factor(s) convergence to the true relevant factors up to a rotation. It is well known in the literature on factor models [^5] This feature of inherent unidentifiability has been emphasized in bai2003, kelly among other papers. The normalization imposed in assumption 5 is done to handle this issue., that true underlying factor(s) are not identifiable; we instead estimate a rotated version of the true factors, which preserves their span.
Define $\boldsymbol{G}_{\beta}\equiv\hat{\boldsymbol{\beta}}_{1}^{-1}{\hat{\boldsymbol{\beta}}_{2}}\left[\boldsymbol{\Lambda}_{f}\boldsymbol{\Delta}_{f}^{3}\boldsymbol{\Lambda}_{f}^{\prime}\right]^{-1}\boldsymbol{\Lambda}_{f}\boldsymbol{\Delta}_{f}^{2}$, where
$\hat{\boldsymbol{\beta}}_{1}=\hat{\boldsymbol{F}}_{A}$ and $\hat{\boldsymbol{\beta}}_{2}=\hat{\boldsymbol{F}}_{B}$, and $\boldsymbol{\Delta}_{f}$ is defined in the paragraph preceding Theorem 1.
@@
$\hat{\boldsymbol{\beta}}-\boldsymbol{G}_{\beta}\boldsymbol{\beta}_{f}=\boldsymbol{O}_{p}(\delta_{MT}^{-1}).$
${\boldsymbol{H}_{f}}^{\prime}\boldsymbol{G}_{\beta}=\boldsymbol{I}_{K_{f}}$
This theorem establishes the convergence of the predictive coefficients to a rotation of the true coefficients. Just like in the case of factor(s), true coefficients are not identifiable and we instead estimate their rotation. The rates established in Theorems 1 and 2 differ from the rates established in kelly and the reason is that our definition of rotation matrices $\boldsymbol{H}_{f}$ and $\boldsymbol{G}_{\beta}$ are different from kelly. (See Remark 1).
@@
As highlighted in bai2006 and also emphasized in kelly, the presence of matrices $\boldsymbol{H}_{f}$ and $\boldsymbol{G}_{\beta}$ in Theorem 1 and 2 highlight the fact we are essentially estimating a vector space. These theorems “pertain to the difference between $\left[\hat{\boldsymbol{F}}_{t}/\hat{\boldsymbol{\beta}}\right$] and the space spanned by $\left[\boldsymbol{F}_{t}/\boldsymbol{\beta}\right]$”. The product $\boldsymbol{H}_{f}^{\prime}\boldsymbol{G}_{\beta}$ equals an identity matrix, cancelling the rotations in the estimated coefficients and the factors; thereby consistently estimating direction spanned by $\boldsymbol{\beta}^{\prime}\boldsymbol{F}_{t}$. However, this characteristic is absent in Theorems 5 and 6 of kelly. The matrices $\boldsymbol{H}$ and $\boldsymbol{G}_{\beta}$ as defined in their paper do not necessarily yield a product that equals an identity matrix.
@@
$\hat{y}_{t+h}-\mathbb{E}_{t}y_{t+h}=O_{p}(\delta_{MT}^{-1}),$
where $\mathbb{E}_{t}y_{t+h}=\beta_{0}+\boldsymbol{\beta}^{\prime}\boldsymbol{F}_{t}=\beta_{0}+\boldsymbol{\beta}_{f}^{\prime}\boldsymbol{f}_{t}$.
Combining Theorem 1 and 2, the convergence of $\hat{y}_{t+h}$ follows directly. Our proof, unlike kelly uses the convergence results for the estimated factor(s) and coefficients to obtain this result.
@@
The rates established in Theorem 1, 2 and 3 are different from the result in kelly where the corresponding rates are $O_{p}({T}^{-1/2})$, $O_{p}({T}^{-1/2})$ and $O_{p}({N}^{-1/2})$ [^6] For our case, it should have been $O_{p}({M}^{-1/2})$ as per their theorem since we apply 3PRF to the transformed M-dimesnional space. respectively (see Theorems 4, 5 and 6 in their paper). For Theorem 1 and 2, the difference is explained by a different definition of the rotation matrices in our paper (see Remark 1). For establishing the convergence of $\hat{y}_{t+h}$, their proof follows two steps. First they show that $\hat{y}_{t+h}-\tilde{y}_{t+h}=O_{p}({T}^{-1/2})$, where $\tilde{y}_{t+h}$ is defined in their appendix. Then then they argue that $\sqrt{T}\tilde{y}_{t+h}\underset{T,N\rightarrow\infty}{\longrightarrow}\mathbb{E}_{t}y_{t+h}$. Since $\tilde{y}_{t+h}$ [^7] The definition of $\tilde{y}_{t+h}$ and fact that it is $O_{p}(1)$ can be seen from the proof of Theorem 4 of kelly is $O_{p}(1)$, $\sqrt{T}\tilde{y}_{t+h}$ would diverge to infinity and their statement would be false. We presume that they erroneously wrote this and instead wanted to imply that $\sqrt{T}\left(\tilde{y}_{t+h}-\mathbb{E}_{t}y_{t+h}\right)\underset{T,N\rightarrow\infty}{\longrightarrow}0$. However this statement is false because $\tilde{y}_{t+h}-\mathbb{E}_{t}y_{t+h}$ has random elements which converge to 0 at a rate which is $O_{p}(M^{-1/2})+O_{p}(T^{-1/2})=O_{p}(\delta_{MT}^{-1}).$